# Efficient 3D Molecular Generation with Flow Matching and Scale Optimal Transport

### v1

@@ Appendix B Ablation Experiments
In Table 4 we present the results of ablation experiments showing the importance of various model components. The results in the table work in the following way – Baseline refers to a MolFlow model with one coordinate set, using equivariant optimal transport and no self conditioning. The models below then build on each other, starting at the top and working down the table, adding the component given in the model column. For example, the model in the row marked by + Scale OT uses 64 coordinate sets and scale optimal transport. The final model is the same as presented in the main text but trained for only 100 epochs, as with the other models in Table 4.
@@
Table 4: MolFlow ablation results on the Geom Drugs dataset. Each model was trained for 100 epochs. During evaluation an average over 5 evaluation runs was taken where each run consisted of sampling 5000 molecules with sizes sampled randomly from the test set.
@@ Appendix C Training Details
All MolFlow models presented in this paper are constructed from 12 Semla layers, with $D=512$, $M=128$ and $S=64$, where $D$ is the model dimension, $M$ is the message dimension, and $S$ is the number of coordinate sets. This corresponds to approximately 40M learnable parameters, which is more than 3 times as many than the current state-of-the-art, EQGAT-diff [26]. Despite this additional expressivity, our model processes batches more than twice as quickly due to the efficiency improvements from Semla.
@@ Appendix D Issues with Current Evaluation Metrics
While conducting evaluations on MolFlow and existing models we have noticed a number of problems with these metrics, which we outline as follows:
@@
Appendix E Samples from MolFlow
In this section we present samples from MolFlow trained on GEOM Drugs. The samples were generated randomly but we have rotated them where necessary to aid visualisation.
Figure 4: Random samples from a MolFlow model trained on GEOM Drugs. These samples were generated using 100 ODE integration steps.

### v2

@@ Appendix B Ablation Experiments
In Table 4 we present the results of ablation experiments showing the importance of various model components. The results in the table work in the following way – Baseline refers to a SemlaFlow model with one coordinate set, using equivariant optimal transport and no self conditioning. The models below then build on each other, starting at the top and working down the table, adding the component given in the model column. For example, the model in the row marked by + Scale OT uses 64 coordinate sets and scale optimal transport. The final model is the same as presented in the main text but trained for only 100 epochs, as with the other models in Table 4.
@@
Table 4: SemlaFlow ablation results on the Geom Drugs dataset. Each model was trained for 100 epochs. During evaluation an average over 5 evaluation runs was taken where each run consisted of sampling 5000 molecules with sizes sampled randomly from the test set.
@@ Appendix C Training Details
All SemlaFlow models presented in this paper are constructed from 12 Semla layers, with $D=512$, $M=128$ and $S=64$, where $D$ is the model dimension, $M$ is the message dimension, and $S$ is the number of coordinate sets. This corresponds to approximately 40M learnable parameters, which is more than 3 times as many than the current state-of-the-art, EQGAT-diff [26]. Despite this additional expressivity, our model processes batches more than twice as quickly due to the efficiency improvements from Semla.
@@ Appendix D Issues with Current Evaluation Metrics
While conducting evaluations on SemlaFlow and existing models we have noticed a number of problems with these metrics, which we outline as follows:
@@
Appendix E Samples from SemlaFlow
In this section we present samples from SemlaFlow trained on GEOM Drugs. The samples were generated randomly but we have rotated them where necessary to aid visualisation.
Figure 4: Random samples from a SemlaFlow model trained on GEOM Drugs. These samples were generated using 100 ODE integration steps.